OpenStudy (anonymous):

help medals will be give !!! At a local plant sale, Janice was selling Golden Barrel Cactus plants and baby Saguaros. During the sale, she sold 5 times as many Saguaros ($6 each) as Golden Barrels ($10 each). She made a total of $720 during the plant sale. How many Saguaros did Janice sell? A. 18 B. 40 C. 72 D. 90

OpenStudy (anonymous):

D.) 90

OpenStudy (anonymous):

5*x*$6 + x*$10 = $720 then if you solve for x you can get the number of golden barrels; Saguaros is five times that.

OpenStudy (anonymous):

is still kinda confusing

OpenStudy (anonymous):

Let x be the number of Golden Barrels sold. IF each GB sells for $10, then 10x is the amount of money made from selling GB. If 5 times as many Saguaros were sold, then 5x is the number of Saguaros sold. Multiply this by $6/Saguaro and you have 30x. 30x is the amount of money made from selling Saguaros. Now add 30x + 10x = 40x. 40x is the total sales from both plants. 40x = 720 x = 18 She sold 5x Saguaros, so 5x = 5*18 = 90

OpenStudy (anonymous):

you can't sell less than one golden barrel cactusses. so, the possible sales from that day are: 5 Saguaros + 1 barrel cactus 10 Saguaros + 2 barrel cactus, 15 Saguaros + 3 barrel cactus, and so on. since the price for each of them is known, it's possible to tell which one of those were the actual numbers.

OpenStudy (anonymous):

5 Saguaros + 1 barrel cactus 10 Saguaros + 2 barrel cactus, 15 Saguaros + 3 barrel cactus x=(1),(2),(3): 5 * (1) $6 + 1*(1) $10 = $40 5 * (2) $6 + 1*(2) $10 = $80 5 * (3) $6 + 1*(3) $10 = $120, and so on 5 * (x) $6 + 1*(x) $10 = $720 you can solve for x to get the actual sell :)

OpenStudy (anonymous):

ohh okayy i get it more thanks u :)

OpenStudy (anonymous):

you're welcome